OpenStudy (anonymous):

Solve for x. 5x - 11 = 7x - 3

OpenStudy (mrnood):

add 11 to each side of the equation THEN subtract 7x from each side of that equaiton

OpenStudy (anonymous):

we want to get all the x terms on one side of the equal sign and everything else on the other side. But remember, what you do on one side, you have to do on the other side. 5x - 11 = 7x - 3 --- subtract 5x from both sides 5x - 5x - 11 = 7x - 5x - 3 -- combine terms -11 = 2x - 3 -- add 3 to both sides -11 + 3 = 2x - 3 + 3 -- combine terms -8 = 2x -- divide both sides by 2 -8/2 = x -- reduce -4 = x check.. 5(-4) - 11 = 7(-4) - 3 -20 - 11 = -28 - 3 -31 = -31 (correct) so x = -4 questions ?